OpenStudy (anonymous):

2 over 3 (x − 4) = 2x for x.

HanAkoSolo (jamierox4ev3r):

If I am understand correctly @Tayloresc , the problem is \(\Large\frac{2}{3(x-4)}=2x\). Is this correct?

OpenStudy (anonymous):

2 over 3 (x-4)=2x for x

HanAkoSolo (jamierox4ev3r):

so is what i wrote correct? lol :P

OpenStudy (anonymous):

no

OpenStudy (anonymous):